OpenStudy (anonymous):

replace each ? with a number to make the statement true... with solutions tnx! (?)^5/3= -32

OpenStudy (saifoo.khan):

Suppose that number be x. x^5/3 = -32 Multiply both sides by exponents 3/5 to cancel the exponents on left.

mathslover (mathslover):

let ? be x : hence \[\large{x^{\frac{5}{3}}=-32}\]

mathslover (mathslover):

-32 = === > \(\large{-8^{\frac{5}{3}}}\)

mathslover (mathslover):

hence : \[\large{x^{\frac{5}{3}}=-8^{\frac{5}{3}}}\] \[\large{x = -8}\] .. m i right ?

OpenStudy (anonymous):

t.y. !!!!!!!!!!!

mathslover (mathslover):

by the way : \[\Huge{\mathbb{Welcome}\textbf{To}\mathbb{Open}\textbf{Study}}\]